Solve using the five-step problem-solving process.
Find the length of a rectangular lot with a perimeter of 108 meters if the length is 8 meters more than the width. (P = 2L + 2W)
thanks

Answer by nyc_function(2741)

length = x + 8
width = x
P = 108
108 = 2(x + 8) + 2(x)
Can you finish now?
=====================
I got your reply. Here is the rest of my reply.
length = x + 8
width = x
P = 108
108 = 2(x + 8) + 2(x)

We have the tools needed to solve for x and find the answer.

Keep in mind that we are searching for the length, which is 8 more than the width denoted as x + 8.

108 = 2(x + 8) + 2(x)

108 = 2x + 16 + 2x

108 = 4x + 16

108 - 16 = 4x

92 = 4x

92/4 = x

23 = x

Is the anwer 23?

No, it is not the answer. The number 23 is the value of x.

Since x = the width, we now know that the width is 23 meters.

Now, the length is x + 8 meters.

Replacing x with 23, we get 23 + 8 or 31 meters.

The length is 31 meters for this rectangular field.